# Transmission Efficiency of the Recoil Mass Spectrometer EMMA at TRIUMF

### v1

@@ 2 Method
respectively, where the incident particles have kinetic energy $T$, mass $m$, and charge $q$ and the spectrometer is set for a reference particle with kinetic energy $T_{0}$, mass $m_{0}$, and charge $q_{0}$. Charged particles entering the spectrometer on the optic axis with $T=T_{0}$, $m=m_{0}$ and $q=q_{0}$ continue to travel along the optic axis. The angles with respect to the optic axis in the horizontal and vertical directions are denoted $\theta$ and $\phi$, respectively.
@@
Therefore, the trajectories of ions in the spectrometer depend only on their $m/q$ and $T/q$ ratios with respect to ions that traverse the spectrometer on the reference trajectory along the optic axis. Similar arguments apply to ions on off-axis trajectories. The transport efficiency, $\epsilon$, is then a function dependent only on the angles $\theta$ and $\phi$ and the ratios $(T/q)/(T_{0}/q_{0})$ and $(m/q)/(m_{0}/q_{0})$, which can be more conveniently expressed via Equation 1 in terms of $\delta T$ and $\delta m$. I.e.,
@@ 3 Measurement Procedure
For these measurements we utilized a 1.18(1) kBq 148Gd $\alpha$ source mounted at the target position. This radionuclide decays 100% of the time via 3.183 MeV $\alpha$ particles [8]. Employing a radioactive source rather than elastically scattered beam or target ions is advantageous for two main reasons. First, all of the $\alpha$ particles emitted by the source have the same charge of $2e$, whereas the charge state distributions of elastically scattered beam and target ions are not precisely known, thereby requiring additional measurements. Second, source measurements do not require an accelerator and can therefore be performed over the extended times required for large $|\delta T|$ settings without preventing accelerated beam experiments at other facilities. The entrance angles of the $\alpha$ particles were defined by apertures fixed within the target chamber 80 mm downstream of the source. The measured aperture dimensions and the horizontal and vertical angles $\theta$ and $\phi$ subtended by each of the 6 apertures are specified in Table 1, along with their solid angles. There is one large aperture, labelled “Full”, which encompasses the entire angular range subtended by each of five smaller apertures, labelled as: “Central”, “Left”, “Right”, “Top” and “Bottom”; the last four apertures are centred off-axis. The orientations and relative sizes of the apertures are depicted in Fig. 2.
@@ 6 Empirical Description
The mean transmission efficiencies calculated according to the best fitting empirical model at each of 9 measured $\delta T$ settings for the 6 angular apertures are shown along with the measurements in Figures 6, 7, 8, 9, 11, and 10. No single model provided the best fit to the set of 6 measurements at every $\delta T$ setting; the offset piecewise Gaussian was the best model for $\delta T>0$, while the piecewise Gaussian was preferred for the $\delta T\leq 0$ measurements except for the two smallest $\delta T$ settings, which required more parameters to obtain acceptable fits. The difference in the quality of the Gaussian and offset Gaussian models fits of the $\delta T=0$ measurement is insignificant. The parameters of the best fitting models for 9 $\delta T$ settings within the interval [0.2,0.2][-0.2,0.2] are shown in Table 3; their $\chi^{2}$ values for the 6 measurements appear in Table 4.
@@
Figure 12 depicts contour plots of the transmission efficiency for the $\delta T=0$ case, calculated according to a piecewise Gaussian function and an offset piecewise Gaussian function, both of whose parameters best fit the 6 angular aperture measurements for $\delta T=0$, as a function of $\theta$ and $\phi$, plotted over the Full angular aperture. The equivalent plots for the $\delta T=-0.1$, $\delta T=-0.05$, and $\delta T=0.05$ cases are shown in Figures 13, 14, 15, and 16, respectively.
@@
The first is the relative uncertainty of the mean transmission efficiency measurements with the Full angular aperture, which includes both statistical and systematic contributions and ranges from 4–6% at the various $\delta T$ settings. The second term is the relative systematic modelling uncertainty. We estimate its size using the difference between the best fitting models for each value of $\delta T$, $\theta$, and $\phi$. E.g., Fig. 17 shows the difference between the transmission efficiency for $\delta T=-0.05$ calculated according to the best-fitting offset and piecewise Gaussian models as a function of $\theta$ and $\phi$, plotted over the Full angular aperture; the average difference between the transmission efficiencies calculated with the two $\delta T=-0.05$ models is 0.012 and the root mean square difference is 0.074 over the angular range of $\pm 3^{\circ}$ by $\pm 3^{\circ}$. The relative systematic modelling uncertainty is taken to be 1/3 of the absolute value of the relative difference between the best fitting models, averaged over the Full angular aperture, and ranges from a minimum of 5% at $\delta T=0$ to a local maximum of 32% at $\delta T=0.2$; an exception is $\delta T=0.2$, where no good fit was obtained. In this case we assigned a 100% relative modelling uncertainty. Table 4 lists the relative measurement and modelling uncertainties for the adopted transmission efficiency models at each of the specified measured $\delta T$ values.
@@
The final term in the quadratic sum is the relative error due to extrapolation beyond the Full angular aperture and accordingly is only finite for trajectories with $|\theta|>3^{\circ}$ or $|\phi|>3^{\circ}$. The absolute uncertainty in this quantity is taken to be 1/3 of the difference between the extrapolated best fitting model and 0. Since the transmission efficiency is assumed to be the average of the two, the relative uncertainty due to extrapolation beyond the Full angular aperture is 2/3.

### v2

@@ 2 Method
respectively, where the incident particles have kinetic energy $T$, mass $m$, and charge $q$ and the spectrometer is set for a reference particle with kinetic energy $T_{0}$, mass $m_{0}$, and charge $q_{0}$. Charged particles entering the spectrometer on the optic axis with $T=T_{0}$, $m=m_{0}$ and $q=q_{0}$ continue to travel along the optic axis. The angles with respect to the optic axis in the horizontal and vertical directions are denoted by $\theta$ and $\phi$, respectively.
@@
Therefore, the trajectories of ions in the spectrometer depend only on their $m/q$ and $T/q$ ratios with respect to ions that traverse the spectrometer on the reference trajectory along the optic axis. Similar arguments apply to ions on off-axis trajectories. The transport efficiency, $\epsilon$, is then a function dependent only on the angles $\theta$ and $\phi$ and the ratios $(T/q)/(T_{0}/q_{0})$ and $(m/q)/(m_{0}/q_{0})$, which can be more conveniently expressed via Equation 1 in terms of $\delta T$ and $\delta m$. That is,
@@ 3 Measurement Procedure
For these measurements we utilized a 1.18(1) kBq 148Gd $\alpha$ source ($T_{1/2}=71.1\pm 1.2$ y) mounted at the target position. This radionuclide decays 100% of the time via 3.183 MeV $\alpha$ particles [8, 9]. Employing a radioactive source rather than elastically scattered beam or target ions is advantageous for two main reasons. First, all of the $\alpha$ particles emitted by the source have the same charge of $2e$, whereas the charge state distributions of elastically scattered beam and target ions are not precisely known, thereby requiring additional measurements. Second, source measurements do not require an accelerator and can therefore be performed over the extended times required for large $|\delta T|$ settings without preventing accelerated beam experiments at other facilities. The entrance angles of the $\alpha$ particles were defined by apertures fixed within the target chamber 80 mm downstream of the source. The measured aperture dimensions and the horizontal and vertical angles $\theta$ and $\phi$ subtended by each of the 6 apertures are specified in Table 1, along with their solid angles. There is one large aperture, labelled “Full”, which encompasses the entire angular range subtended by each of five smaller apertures, labelled as: “Central”, “Left”, “Right”, “Top” and “Bottom”; the last four apertures are centred off-axis. The orientations and relative sizes of the apertures are depicted in Fig. 2.
@@ 6 Empirical Description
The mean transmission efficiencies calculated according to the best fitting empirical model at each of 9 measured $\delta T$ settings for the 6 angular apertures are shown along with the measurements in Figures 6, 7, 8, 9, 10, and 11. No single model provided the best fit to the set of 6 measurements at every $\delta T$ setting; the offset piecewise Gaussian was the best model for $\delta T>0$, while the piecewise Gaussian was preferred for the $\delta T\leq 0$ measurements except for the two smallest $\delta T$ settings, which required more parameters to obtain acceptable fits. The difference in the quality of the Gaussian and offset Gaussian model fits of the $\delta T=0$ measurements is insignificant. The parameters of the best fitting models for 9 $\delta T$ settings within the interval [0.2,0.2][-0.2,0.2] are shown in Table 3; their $\chi^{2}$ values for the 6 measurements appear in Table 4.
@@
Figure 12 depicts contour plots of the transmission efficiency for the $\delta T=0$ case, calculated according to a piecewise Gaussian function and an offset piecewise Gaussian function, both of whose parameters best fit the 6 angular aperture measurements for $\delta T=0$, as a function of $\theta$ and $\phi$, plotted over the Full angular aperture. The equivalent plots for the $\delta T=-0.1$, $\delta T=-0.05$, $\delta T=0.05$, and $\delta T=0.1$ cases are shown in Figures 13, 14, 15, and 16, respectively.
@@
The first is the relative uncertainty of the mean transmission efficiency measurements with the Full angular aperture, which includes both statistical and systematic contributions and ranges from 4–6% at the various $\delta T$ settings. The second term is the relative systematic modelling uncertainty. We estimate its size using the difference between the best fitting models for each value of $\delta T$, $\theta$, and $\phi$. For example, Fig. 17 shows the difference between the transmission efficiency for $\delta T=-0.05$ calculated according to the best-fitting offset and piecewise Gaussian models as a function of $\theta$ and $\phi$, plotted over the Full angular aperture; the average difference between the transmission efficiencies calculated with the two $\delta T=-0.05$ models is 0.012 and the root mean square difference is 0.074 over the angular range of $\pm 3^{\circ}$ by $\pm 3^{\circ}$. The relative systematic modelling uncertainty is taken to be 1/3 of the absolute value of the relative difference between the best fitting models, averaged over the Full angular aperture, and ranges from a minimum of 5% at $\delta T=0$ to a local maximum of 32% at $\delta T=0.2$; an exception is $\delta T=-0.2$, where no good fit was obtained. In this case we assigned a 100% relative modelling uncertainty. Table 4 lists the relative measurement and modelling uncertainties for the adopted transmission efficiency models at each of the specified measured $\delta T$ values.
@@
The final term in the quadratic sum is the relative error due to extrapolation beyond the Full angular aperture and accordingly is only finite for trajectories with $|\theta|>3^{\circ}$ or $|\phi|>3^{\circ}$ (0.052 rad). The absolute uncertainty in this quantity is taken to be 1/3 of the difference between the extrapolated best fitting model and 0. Since the transmission efficiency is assumed to be the average of the two, the relative uncertainty due to extrapolation beyond the Full angular aperture is 2/3.